OpenStudy (anonymous):

solve for Y. if 23x78=y

OpenStudy (anonymous):

would i have to divid?

OpenStudy (superhelp101):

multiply

OpenStudy (anonymous):

its 1794

OpenStudy (anonymous):

but wouldn't i have to do the opposite of multiplication for this problem?

OpenStudy (superhelp101):

why though. it says 23X78=Y so u just have to find Y by multiplying those two numbers :)

OpenStudy (anonymous):

yep!

OpenStudy (superhelp101):

YaY

OpenStudy (anonymous):

no, all u do is multiple

OpenStudy (superhelp101):

@cucu u got it?

OpenStudy (anonymous):

my teacher told me a whole diff story lol

OpenStudy (anonymous):

like for an example t+28=54 i would have to subtract 28 both sides

OpenStudy (anonymous):

haha wow, its simple cause all u do is multiply

OpenStudy (anonymous):

there no other operation sign except the multiply

OpenStudy (superhelp101):

but it says Y is equal to 23X78 so if Y is equal to that then u just have to multi. the two number to get the answer :D

OpenStudy (anonymous):

ohhh so how would i show my work

OpenStudy (superhelp101):

and yes @cucu u are correct you have to subtract. t+28=54 -28 -28 t=26

OpenStudy (anonymous):

noo I'm talking about the other problem 28x78=y??

OpenStudy (anonymous):

how would that work be shown?

OpenStudy (superhelp101):

do you know how to multi ? :)

OpenStudy (anonymous):

yeah i got 1794

OpenStudy (anonymous):

how would you check your answer ?

OpenStudy (superhelp101):

well that is what Y equals so u you check nothing :D

OpenStudy (anonymous):

lol ok thanks for the help

OpenStudy (superhelp101):

your welcome ;P

OpenStudy (anonymous):

your very welcome! let me know if u have any other questions